OpenStudy (anonymous):

Please help me solve this A numbers ones digit is 2 more than twice its tens digit. The sum of its digits is 5. What's the number?

OpenStudy (anonymous):

Let's say the number is ab, a is the tens digit and b is the ones digit. 2a+2=b and a+b=5 Substitute b a+2a+2=5 3a+2=5 a=1 b=4 14